OpenStudy (anonymous):

what is -2sqrt2p*2sqrt22

OpenStudy (anonymous):

\[-2\sqrt{2p}*2\sqrt{22}\]

OpenStudy (phi):

First multiply the two numbers not under the square roots. what do you get?

OpenStudy (anonymous):

44p

OpenStudy (phi):

only the -2 and 2 are NOT under the square root. what do you get?

OpenStudy (anonymous):

-4

OpenStudy (phi):

good. now look at the numbers under the square root. We have \[ \sqrt{2}\cdot\sqrt{22}\cdot\sqrt{p} \] If we split them. We can combine the numbers: \[ \sqrt{2 \cdot 22}\cdot \sqrt{p} \]

OpenStudy (phi):

Now the important part. Notice that 22= 2*11, so we have under the square root 2*2*11 and the 2*2 is a square. Putting it all together (remember the -4 ) \[ −4\cdot 2\sqrt{11}\cdot\sqrt{p} \] or \[ −8\sqrt{11p} \]